# “6 choose 4”: A framework to understand and facilitate discussion of strategies for overall survival safety monitoring

### v2

@@ Abstract
Advances in anticancer therapies have contributed significantly to decreasing death rates in certain disease and clinical settings. However, they have also made it difficult to power a clinical trial in these settings with overall survival (OS) as the primary efficacy endpoint. Therefore, two approaches have recently been proposed for the pre-specified analysis of OS as a safety endpoint (Fleming et al., 2024; Rodriguez et al., 2024). We provide in this paper a simple unifying framework that includes the aforementioned approaches—and a few others—as special cases. By highlighting each approach’s focus, priority, tolerance for risk, and strengths or challenges for practical implementation, this framework can help facilitate discussions between stakeholders on “fit-for-purpose OS data collection and assessment of harm” (American Association for Cancer Research, 2023). We apply this framework to a real clinical trial in large B-cell lymphoma to illustrate its application and value. Several recommendations and open questions are also raised.
@@ 1 Introduction
With the continued advancement of anticancer therapies, there are now many disease and clinical contexts in which patients’ lives are extended and multiple treatment options are available to patients following progression. Although these successes should be celebrated, they have also made it challenging for clinical trials to have adequate power to detect statistically significant improvements in overall survival (OS)—the gold standard endpoint in oncology (Merino et al., 2024). In July 2023, a workshop titled “Overall Survival in Oncology Clinical Trials” was held jointly by the U.S. Food and Drug Administration (FDA), American Association for Cancer Research (AACR), and American Statistical Association (ASA) to discuss challenges with the timely evaluation of OS for novel therapies (American Association for Cancer Research, 2023). The workshop cited lack of OS data, lack of plans for further OS data collection, and lack of pre-specified OS analyses as emerging challenges for regulatory authorities to interpret OS data in trials where OS is not a primary or key secondary endpoint. A call to action was made to “encourage thoughtful and comprehensive planning during trial design for fit-for-purpose OS data collection and assessment of harm given the disease and clinical context”.
@@
Following that call-to-action, two statistical approaches were proposed for the pre-specified analysis of OS as a safety endpoint. Fleming et al. (2024) proposed a monitoring guideline inspired by ones used by the FDA in cardiovascular safety trials. Rodriguez et al. (2024) suggested using confidence intervals to decide whether or not there is sufficient evidence to rule out harm. These approaches are similar to a certain degree. For example, both consider similar sets of parameters: definition of harm measured in terms of the OS hazard ratio, type I error, etc.. However, upon closer look, it can be unclear even to the experienced statistician how these approaches relate to each other exactly, how they differ, which approach one should use in practice, and why.
@@
This paper proposes a framework for OS safety analysis that includes as special cases Fleming et al. (2024) and Rodriguez et al. (2024). Importantly, this framework explains what flexibility and restrictions exist for any approach analyzing OS as a safety endpoint under the traditional framework of hypothesis testing, and clarifies for each approach their focus, priority, and tolerance for risk over the course of a trial. By increasing understanding of different approaches and the relationship between them, this framework can help facilitate “thoughtful planning” by sponsors and improve communication between stakeholders in order to achieve alignment on clinical trial design and analysis.
@@
The rest of the paper is organized as follows. In Section 2, we present our framework for OS safety monitoring and provide five examples of approaches that fall under this framework, including Fleming et al. (2024) and Rodriguez et al. (2024). In Section 3, we illustrate and compare the five approaches by applying them to POLARIX, a real clinical trial studying patients with untreated large B-cell lymphoma. Section 4 concludes with recommendations for practice and open questions for further deliberation or research.
@@ 2 Proposed framework
For the purpose of safety monitoring in indolent cancer trials, we may limit ourselves to considering $\theta_{0}>1.0$ and moderate to neutral $\theta_{1}$ (e.g., between 0.7 and 1.0). Fleming et al. (2024) characterized $\theta_{0}$ as “the smallest unacceptable detrimental OS HR” and $\theta_{1}$ as “a plausible OS HR consistent with benefit that is reasonably expected from the experimental intervention”.
@@
Given six parameters and Equations (I)-(II) characterizing the relationship between them, our proposed framework is as follows: at each analysis time, users input or “choose” 4 of the 6 parameters (not all from the same equation) and solve for the remaining 2. This one-sentence framework communicates clearly what flexibility and restrictions exist in any OS monitoring strategy. It highlights a strategy’s focus and associated risks; for example, choosing $\beta$ and solving for $\alpha$ at an early interim analysis implies a focus on having the trial continue in case the experimental therapy is effective, while understanding the risk of a false positive. And as we shall illustrate, it includes as special cases Fleming et al. (2024), Rodriguez et al. (2024), and other approaches for OS safety monitoring.
@@
Rodriguez et al. proposed to calculate the $100\times(1-2\alpha)\%$ confidence interval (CI) for the estimated OS HR, at a time when OS data is relatively mature. Failure to exclude $\theta_{0}$ with the upper confidence bound is then seen as an indication of OS detriment. Their approach is equivalent to choosing a single $(\theta_{0},\theta_{1},\alpha,\beta)$ and solving for $(d,\theta^{*})$ (Shan, 2023). In other words, like traditional hypothesis testing and sample size calculation, Rodriguez et al. use a CI for decision making precisely at a time when a desirable set of operating characteristics can be achieved.
@@
An alternative approach to Rodriguez et al. (2024) that one might consider is to monitor OS safety using $100\times(1-2\alpha)\%$ CIs again but with flexibility for number of events, i.e., by choosing $(\theta_{0},\theta_{1},\alpha,d)$ at each analysis time. (The choice between $\theta_{1}$ or $\beta$ is flexible and depends on the risk one would like to quantify.)
An example of this approach is an independent data monitoring committee (IDMC) who may not have been provided with detailed instructions for OS safety monitoring and is continually using 95% CIs for evaluation. As commented by Dr. Mikkael Sekeres at the FDA-AACR-ASA workshop, “It is surprising how little guidance you’re often given when you’re on an IDMC… IDMC should be given good guidance on how to consider overall survival… We think that IDMC should be given specific tasks to look at overall survival at different stages of a trial and also statistical guidance on the confidence with which you can declare that a drug is either too harmful or is safe in moving forward” (American Association for Cancer Research, 2023).
@@
Our final example is to choose $(\theta_{1},\alpha,\beta,d)$ at IAs and $(\theta_{0},\theta_{1},\alpha,d)$ at the FA. This approach is based on monitoring guidelines that the FDA has used in trials of type 2 diabetes mellitus (T2D) to evaluate cardiovascular risk (FDA Center for Drug Evaluation and Research, 2008), guidelines which also served to inspire approach #1 by Fleming et al.. The key difference between approach #5 and #1 is that, while approach #1 addresses the question “With what certainty can we rule out a particular level of OS detriment?”, approach #5 places more emphasis on type I error control at IAs and addresses the question “What OS detriment can we rule out with a high level of certainty?”.
@@ 3 Case study
POLARIX was a Phase 3 randomized controlled trial that compared Pola-R-CHP and R-CHOP in patients with untreated large B-cell lymphoma (LBCL). At the time of the primary analysis for PFS and first IA for OS, the trial provided statistically significant evidence of PFS benefit (PFS HR estimate 0.73; 95% CI 0.57 to 0.95), but immature and inconclusive OS data despite all patients having a minimum of 2 years follow-up (OS HR estimate 0.94, 95% CI 0.67 to 1.33). Subsequently, an FDA ODAC meeting was held to discuss the benefit-risk of Pola-R-CHP, taking into consideration the uncertainty around OS due to low number of deaths. For a detailed account of POLARIX and the ODAC’s eventual vote in favor of Pola-R-CHP, we refer the reader to the briefing document (FDA ODAC, 2023).
Fleming et al. (2024) illustrated their OS monitoring guideline by retrospectively applying it to POLARIX (Strategy 1, Table 1). Under our proposed framework, their guideline is equivalent to choosing $(d,\theta_{0},\theta_{1},\beta)=(89,1.30,0.80,0.1)$ at IA1, $(d,\theta_{0},\theta_{1},\beta)=(110,1.30,0.80,0.1)$ at IA2, $(d,\theta_{0},\theta_{1},\beta)=(131,1.30,0.80,0.1)$ at IA3, and $(d,\theta_{0},\theta_{1},\alpha)=(178,1.30,0.80,0.025)$ at FA. For the rationale behind some of these parameters, we refer the reader to Fleming et al. (2024). We now illustrate and compare the remaining four approaches described in Section 2 by choosing similar parameter values wherever the set of input parameters overlap with Fleming et al.’s, and choosing conventional/intuitive values wherever they do not.
@@
3A. Standard IA1 89 1.30 0.80 0.858 0.025 0.629
CIs IA2 110 1.30 0.80 0.895 0.025 0.721
IA3 131 1.30 0.80 0.923 0.025 0.793
FA 178 1.30 0.80 0.969 0.025 0.900
Overall 0.006a 0.584b
Table 1 (continued)
3B. Standard IA1 89 1.30 0.80 1.044 0.150 0.895
CIs IA2 110 1.30 0.80 1.018 0.100 0.897
@@
If there is a desire to control both type I and type II error rates, then one approach is to follow Rodriguez et al. (2024) by choosing $(\theta_{0},\theta_{1},\alpha,\beta)$ over time. Strategy 2 in Table 1 specifies standard type I error rates that decrease over time (0.05 and 0.025) and type II error rate 0.10. The resulting thresholds $\theta^{*}$ are similar to those of Strategy 1 at IA3 and FA. However, the actual timing of IA in POLARIX was driven by PFS, not OS; the trial targeted 228 PFS events, eventually observing 241 PFS events along with 110 OS events. Should POLARIX have assessed for potential OS detriment then, and if so, how? Or should it have waited until 145 and 178 OS events before assessing for potential OS detriment? The latter may seem odd considering that in reality POLARIX formally assessed OS for efficacy at IA under the group sequential framework.
Strategy 3A gives back flexibility on the timing of safety monitoring by continually using 95% CIs to rule out harm. It is equivalent to choosing $(d,\theta_{0},\theta_{1},\alpha=0.025)$ at all times. Expecting 95% CIs to rule out $\theta_{0}$ early on in POLARIX results in very stringent thresholds (0.858 with 89 deaths, 895 with 110 deaths, and 0.923 with 131 deaths) and substantially lower probabilities of claiming OS safety when Pola-R-CHP is in fact effective (0.629, 0.721, and 0.793). It is for this reason that we do not recommend stringent control of type I error early on in a trial, since immature OS data and wide 95% CIs can cause unnecessary alarm and result in the wrongful termination of effective treatments.
Compared to Strategy 3A, Strategy 3B uses 70%, 80%, and 90% CIs to rule out harm early on in POLARIX. It is equivalent to choosing $(d=89,\theta_{0},\theta_{1},\alpha=0.15)$, $(d=110,\theta_{0},\theta_{1},\alpha=0.1)$, $(d=131,\theta_{0},\theta_{1},\alpha=0.05)$, and $(d=178,\theta_{0},\theta_{1},\alpha=0.025)$. Strategy 3B results in more relaxed thresholds and higher probabilities of claiming OS safety when Pola-R-CHP is effective, similar to those of Strategy 1.
@@
Besides the marginal risks involved at each stage in a trial, stakeholders should also consider the overall risk of the trial should the decision making thresholds $\theta^{*}$ be strictly followed. Table 1 calculates for each strategy the probability of consistently being deemed safe (i.e., probability of meeting all thresholds $\theta^{*}$) under $H_{0}$ and under $H_{1}$. It can be seen that all strategies have a high probability $(>97\%)$ of being flagged as unsafe at least once (i.e., failing to meet a threshold at least once) during trial conduct if the drug is harmful. Meanwhile, five of the strategies have a >>80% probability of consistently being deemed safe if the drug is effective, the only exception being Strategy 3A (58%).
For Strategy 3A, the family wise error rate (i.e., probability of being deemed safe at least once under $H_{0}$) is 0.053 (not shown in Table 1). If there is a desire to control this rate at a lower level, say 0.05 or 0.025, then one can expect a less than 58% probability of consistently being deemed safe under $H_{1}$—a level of risk that sponsors might be unwilling to take. In the context of safety monitoring where interest is in flagging a treatment that causes OS detriment, we think the probability of consistently being deemed safe under $H_{0}$—or more precisely, its complement, the probability of being flagged as unsafe at least once if the drug is harmful—is a more meaningful quantity than the FWER.
@@ 4 Discussion
We described five special cases under the proposed framework: Fleming et al. (2024), Rodriguez et al. (2024), an approach that uses standard confidence intervals to continually rule out harm, another approach that uses discrete thresholds for decision making, and a fifth approach that is based off of previous FDA monitoring guidelines for evaluating cardiovascular risk in T2D. We illustrated use of the five approaches by applying them to the POLARIX trial (Section 3). While the main objective of our paper is to increase understanding of each approach, thereby facilitating planning by sponsors of future clinical trials and improving communication between stakeholders, we leave here with some final comments and recommendations regarding their use.
@@
More relaxed CIs (90%, 80%, etc.) or discrete thresholds may be appropriate provided that the risks involved are understood and acceptable. In fact, one could apply Fleming et al. as a first approximation, then find an approach using standard CIs or discrete thresholds that has similar operating characteristics. Such an approach could effectively balance priorities over time and be easy to implement during a trial. Sponsors taking this approach would do well to pre-specify when and how to adjust the $\alpha$-level or discrete threshold in case the number of events deviates significantly from expectations during trial conduct (see Figure 1 for example).
Figure 1: Fleming et al. (2024) and an approximately equivalent discrete threshold strategy applied to POLARIX at the first planned interim analysis with 89 expected OS events. Solid curve is the hazard ratio (HR) threshold for decision making based on Fleming et al. (2024) with $\theta_{0}=1.3$, $\theta_{1}=0.80$, $1-\beta=0.9$, and varying potential number of observed deaths $d$. Dashed curve indicates discrete thresholds that approximate Fleming et al. (2024) by maintaining $1-\beta\geq 0.9$ while tolerating higher levels of $\alpha$.
Regardless of the approach taken, the low mortality rate in indolent diseases makes rigorous OS monitoring challenging. We agree with Fleming et al. (2024) that any set of proposed thresholds should be seen as non-binding “informative guidelines” as opposed to black-and-white “rules” for continuing or stopping a trial. At a minimum, the need to monitor OS for safety should encourage stakeholders to contemplate the level of OS data that is likely to be accessible during the course of a clinical trial. Meanwhile, the additional analyses suggested by Rodriguez et al. (2024) may be helpful for gaining a better understanding of immature OS data: analysis of biologically plausible subgroups, simulation and supplementary analyses to investigate the robustness of results under different assumptions, evaluation of other endpoints, results from other trials, use of real-world data, etc.. IDMCs may also find measures of stability or graphical summaries such as the predicted interval plot useful for understanding potential effect size estimates and associated precision with trial continuation (Betensky, 2015; Li et al., 2009).
@@
  • How should $\theta^{*}$ be defined given potential confounding by intercurrent events? If a substantial number of patients switch to subsequent therapy after tumor progression, then it may be desirable to relax $\theta^{*}$ since subsequent therapies can lead to a smaller OS difference (Korn et al., 2011).
@@
Another question that may come up is regarding non-proportional hazards (NPH). Is the proposed framework applicable under NPH? Our opinion is yes, that at least a good part of the framework can be applied if the overall HR is still considered a meaningful treatment effect summary; for example, analysis timing can still be based on the number of deaths $d$ and $\theta^{*}$ can still be used as the threshold for signaling overall OS detriment. However, if it is known a priori that a NPH effect could exist, then operating characteristics should be evaluated carefully under potential patterns of NPH in order to ensure that whatever plan that’s proposed for OS safety analysis balances risk appropriately over time. Furthermore, supportive analyses such as piecewise exponential and cure rate models can be prospectively planned to help interpret trial results. This is similar to how some sponsors currently handle NPH for efficacy analysis. They conduct event-driven analyses using the logrank test and cox HR, but consider potential NPH when calculating the required number of events. For example, see Checkmate-274 and Keynote-966 (Bajorin et al., 2021; Kelley et al., 2023).
This paper proposed a supportive decision making framework based on a single criteria $\theta^{*}$. It is worth noting that Rodriguez et al. (2024) considered having a ‘grey zone’, which, if a trial were to fall into, would prompt further OS follow-up. Doing so could have the effect of ruling out substantial OS detriment while limiting rejection of treatments with marginal OS benefit. While we did not consider such an extension here for simplicity of comparison and presentation, we do find this idea of practical interest worth pursuing in future discussions and research.

### v3

@@ Abstract
Advances in anticancer therapies have contributed significantly to decreasing death rates in certain disease and clinical settings. However, they have also made it difficult to power a clinical trial in these settings with overall survival (OS) as the primary efficacy endpoint. Therefore, two approaches have recently been proposed for the pre-specified analysis of OS as a safety endpoint (fleming; rodriguez). We provide in this paper a simple unifying framework that includes the aforementioned approaches—and a few others—as special cases. By highlighting each approach’s focus, priority, tolerance for risk, and strengths or challenges for practical implementation, this framework can help facilitate discussions between stakeholders on “fit-for-purpose OS data collection and assessment of harm” (aacr). We apply this framework to a real clinical trial in large B-cell lymphoma to illustrate its application and value. Several recommendations and open questions are also raised.
@@ 1 Introduction
With the continued advancement of anticancer therapies, there are now many disease and clinical contexts in which patients’ lives are extended and multiple treatment options are available following disease progression. Although these successes should be celebrated, they have also made it challenging for clinical trials to have adequate power to detect statistically significant improvements in overall survival (OS)—the gold standard endpoint in oncology (Merino2024). In July 2023, a workshop titled “Overall Survival in Oncology Clinical Trials” was held jointly by the U.S. Food and Drug Administration (FDA), American Association for Cancer Research (AACR), and American Statistical Association (ASA) to discuss challenges with the timely evaluation of OS for novel therapies (aacr). The workshop cited lack of OS data, lack of plans for further OS data collection, and lack of pre-specified OS analyses as emerging challenges for regulatory authorities to interpret OS data in trials where OS is not a primary or key secondary endpoint. A call to action was made to “encourage thoughtful and comprehensive planning during trial design for fit-for-purpose OS data collection and assessment of harm given the disease and clinical context”.
@@
Following that call-to-action, two statistical approaches were proposed for the pre-specified analysis of OS as a safety endpoint. fleming proposed a monitoring guideline inspired by ones used by the FDA in cardiovascular safety trials. rodriguez suggested using confidence intervals to decide whether or not there is sufficient evidence to rule out harm. These approaches are similar to a certain degree. For example, both consider similar sets of parameters: definition of harm measured in terms of the OS hazard ratio, type I error, etc.. However, upon closer look, it can be unclear even to the experienced statistician how these approaches relate to each other exactly, how they differ, which approach one should use in practice, and why.
@@
This paper proposes a framework for OS safety analysis that includes as special cases fleming and rodriguez. Importantly, this framework explains what flexibility and restrictions exist for any approach analyzing OS as a safety endpoint under the traditional framework of hypothesis testing, and clarifies for each approach their focus, priority, and tolerance for risk over the course of a trial. By increasing understanding of different approaches and the relationship between them, this framework can help facilitate “thoughtful planning” by sponsors and improve communication between stakeholders in order to achieve alignment on clinical trial design and analysis.
@@
The rest of the paper is organized as follows. In Section 2, we present our framework for OS safety monitoring and provide five examples of approaches that fall under this framework, including fleming and rodriguez. In Section 3, we illustrate and compare the five approaches by applying them to POLARIX, a real clinical trial studying patients with untreated large B-cell lymphoma. Section 4 concludes with recommendations for practice and open questions for further deliberation or research.
@@ 2 Proposed framework
For the purpose of safety monitoring in indolent cancer trials, we may limit ourselves to considering $\theta_{0}>1.0$ and moderate to neutral $\theta_{1}$ (e.g., between 0.7 and 1.0). fleming characterized $\theta_{0}$ as “the smallest unacceptable detrimental OS HR” and $\theta_{1}$ as “a plausible OS HR consistent with benefit that is reasonably expected from the experimental intervention”.
@@
Given six parameters and Equations (I)-(II) characterizing the relationship between them, our proposed framework is as follows: at each analysis time, users input or “choose” 4 of the 6 parameters (not all from the same equation) and solve for the remaining 2. This one-sentence framework communicates clearly what flexibility and restrictions exist in any OS monitoring strategy. It highlights a strategy’s focus and associated risks; for example, choosing $\beta$ and solving for $\alpha$ at an early interim analysis implies a focus on having the trial continue in case the experimental therapy is effective, while understanding the risk of a false positive. And as we shall illustrate, it includes as special cases fleming, rodriguez, and other approaches for OS safety monitoring.
@@
Rodriguez et al. proposed to calculate the $100\times(1-2\alpha)\%$ confidence interval (CI) for the estimated OS HR, at a time when OS data is relatively mature. Failure to exclude $\theta_{0}$ with the upper confidence bound is then seen as an indication of OS detriment. Their approach is equivalent to choosing a single $(\theta_{0},\theta_{1},\alpha,\beta)$ and solving for $(d,\theta^{*})$ (shan). In other words, like traditional hypothesis testing and sample size calculation, Rodriguez et al. use a CI for decision making precisely at a time when a desirable set of operating characteristics can be achieved.
@@
An alternative approach to rodriguez that one might consider is to monitor OS safety using $100\times(1-2\alpha)\%$ CIs again but with flexibility for number of events, i.e., by choosing $(\theta_{0},\theta_{1},\alpha,d)$ at each analysis time. (The choice between $\theta_{1}$ or $\beta$ is flexible and depends on the risk one would like to quantify.)
An example of this approach is an independent data monitoring committee (IDMC) who may not have been provided with detailed instructions for OS safety monitoring and is continually using 95% CIs for evaluation. As commented by Dr. Mikkael Sekeres at the FDA-AACR-ASA workshop, “It is surprising how little guidance you’re often given when you’re on an IDMC… IDMC should be given good guidance on how to consider overall survival… We think that IDMC should be given specific tasks to look at overall survival at different stages of a trial and also statistical guidance on the confidence with which you can declare that a drug is either too harmful or is safe in moving forward” (aacr).
@@
Our final example is to choose $(\theta_{1},\alpha,\beta,d)$ at IAs and $(\theta_{0},\theta_{1},\alpha,d)$ at the FA. This approach is based on monitoring guidelines that the FDA has used in trials of type 2 diabetes mellitus (T2D) to evaluate cardiovascular risk (fda2008), guidelines which also served to inspire approach #1 by Fleming et al.. The key difference between approach #5 and #1 is that, while approach #1 addresses the question “With what certainty can we rule out a particular level of OS detriment?”, approach #5 places more emphasis on type I error control at IAs and addresses the question “What OS detriment can we rule out with a high level of certainty?”.
@@ 3 Case study
POLARIX was a Phase 3 randomized controlled trial that compared Pola-R-CHP and R-CHOP in patients with untreated large B-cell lymphoma (LBCL). At the time of the primary analysis for PFS and first IA for OS, the trial provided statistically significant evidence of PFS benefit (PFS HR estimate 0.73; 95% CI 0.57 to 0.95), but immature and inconclusive OS data despite all patients having a minimum of 2 years follow-up (OS HR estimate 0.94, 95% CI 0.67 to 1.33). Subsequently, an FDA ODAC meeting was held to discuss the benefit-risk of Pola-R-CHP, taking into consideration the uncertainty around OS due to low number of deaths. For a detailed account of POLARIX and the ODAC’s eventual vote in favor of Pola-R-CHP, we refer the reader to the briefing document (fda).
fleming illustrated their OS monitoring guideline by retrospectively applying it to POLARIX (Strategy 1, Table 1). Under our proposed framework, their guideline is equivalent to choosing $(d,\theta_{0},\theta_{1},\beta)=(89,1.30,0.80,0.1)$ at IA1, $(d,\theta_{0},\theta_{1},\beta)=(110,1.30,0.80,0.1)$ at IA2, $(d,\theta_{0},\theta_{1},\beta)=(131,1.30,0.80,0.1)$ at IA3, and $(d,\theta_{0},\theta_{1},\alpha)=(178,1.30,0.80,0.025)$ at FA. For the rationale behind some of these parameters, we refer the reader to fleming. We now illustrate and compare the remaining four approaches described in Section 2 by choosing similar parameter values wherever the set of input parameters overlap with Fleming et al.’s, and choosing conventional/intuitive values wherever they do not.
@@
3. Standard CIs IA1 89 1.30 0.80 1.044 0.150 0.895
IA2 110 1.30 0.80 1.018 0.100 0.897
@@
If there is a desire to control both type I and type II error rates, then one approach is to follow rodriguez by choosing $(\theta_{0},\theta_{1},\alpha,\beta)$ over time. Strategy 2 in Table 1 specifies standard type I error rates that decrease over time (0.05 and 0.025) and type II error rate 0.10. The resulting thresholds $\theta^{*}$ are similar to those of Strategy 1 at IA3 and FA. However, the actual timing of IA in POLARIX was driven by PFS, not OS; the trial targeted 228 PFS events, eventually observing 241 PFS events along with 110 OS events. Should POLARIX have assessed for potential OS detriment then, and if so, how? Or should it have waited until 145 and 178 OS events before assessing for potential OS detriment? The latter may seem odd considering that in reality POLARIX formally assessed OS for efficacy at IA under the group sequential framework.
Strategy 3 gives back flexibility on the time of safety monitoring by using 70%, 80%, and 90% CIs to rule out harm early on in POLARIX. It is equivalent to choosing $(d=89,\theta_{0},\theta_{1},\alpha=0.15)$, $(d=110,\theta_{0},\theta_{1},\alpha=0.1)$, $(d=131,\theta_{0},\theta_{1},\alpha=0.05)$, and $(d=178,\theta_{0},\theta_{1},\alpha=0.025)$. The results are more relaxed thresholds and higher probabilities of claiming OS safety when Pola-R-CHP is effective, similar to those of Strategy 1.
If 95% CIs were to be used instead from IA1 all the way through to FA (not shown in Table 1), then Strategy 3 would result in very stringent thresholds (0.858 with 89 deaths, 895 with 110 deaths, and 0.923 with 131 deaths) and substantially lower probabilities of claiming OS safety when Pola-R-CHP is in fact effective (0.629, 0.721, and 0.793). It is for this reason that we do not recommend stringent control of type I error early on in a trial, since immature OS data and wide 95% CIs can cause unnecessary alarm and result in the wrongful termination of effective treatments.
@@
Besides the marginal risks involved at each stage in a trial, stakeholders should also consider the overall risk of the trial should the decision making thresholds $\theta^{*}$ be strictly followed. Table 1 calculates for each strategy the probability of consistently being deemed safe (i.e., probability of meeting all thresholds $\theta^{*}$) under $H_{0}$ and under $H_{1}$. It can be seen that all strategies have a high probability $(>97\%)$ of being flagged as unsafe at least once (i.e., failing to meet a threshold at least once) during trial conduct if the drug is harmful. They also have a >>80% probability of consistently being deemed safe if the drug is effective.
For Strategies 1, 3, 4, and 5, the family wise error rates (i.e., probability of being deemed safe at least once under $H_{0}$) are between 0.17 to 0.24 (not shown in Table 1). If there is a desire to control these rates at a lower level, say 0.05 or 0.025, then one can expect a less than 80% probability of consistently being deemed safe under $H_{1}$—a level of risk that sponsors might be unwilling to take. In the context of safety monitoring where interest is in flagging a treatment that causes OS detriment, we think the probability of consistently being deemed safe under $H_{0}$—or more precisely, its complement, the probability of being flagged as unsafe at least once if the drug is harmful—is a more meaningful quantity than the FWER.
@@ 4 Discussion
We described five special cases under the proposed framework: fleming, rodriguez, an approach that uses standard confidence intervals to continually rule out harm, another approach that uses discrete thresholds for decision making, and a fifth approach that is based off of previous FDA monitoring guidelines for evaluating cardiovascular risk in T2D. We illustrated use of the five approaches by applying them to the POLARIX trial (Section 3). While the main objective of our paper is to increase understanding of each approach, thereby facilitating planning by sponsors of future clinical trials and improving communication between stakeholders, we leave here with some final comments and recommendations regarding their use.
@@
The other limitation of Rodriguez et al.—the limitation that planned vs. observed number of deaths are typically different—can be addressed by extending the authors’ approach within our proposed framework. At the design stage, sponsors can use Rodriguez et al. to define and predict the number of PFS and OS events at an IA, seeking a balance between properly powering PFS and observing sufficiently mature OS. Subsequently, a threshold $\theta^{*}$ can be calculated based on the designed $d$. At the IA, if the observed number of deaths $d^{\prime}$ is different from the designed $d$, then one can use $(\theta_{0},\theta_{1},\alpha,d^{\prime})$ or $(\theta_{0},\theta_{1},\beta,d^{\prime})$ to update $\theta^{*}$ depending on what risk one could like to quantify. This is similar to the concept of group sequential designs, where stopping boundaries are updated prior to formal readouts based on the observed numbers of deaths.
More relaxed CIs (90%, 80%, etc.) or discrete thresholds may be appropriate provided that the risks involved are understood and acceptable. In fact, one can apply Fleming et al. as a first approximation, then find an approach using standard CIs or discrete thresholds that has similar operating characteristics. Such an approach could effectively balance priorities over time and be easy to implement during a trial. Sponsors taking this approach would do well to pre-specify when and how to adjust the $\alpha$-level or discrete threshold in case the number of events deviates significantly from expectations during trial conduct (see Figure 1 for example).
Figure 1: fleming and an approximately equivalent discrete threshold strategy applied to POLARIX at the first planned interim analysis with 89 expected OS events. Solid curve is the hazard ratio (HR) threshold for decision making based on fleming with $\theta_{0}=1.3$, $\theta_{1}=0.80$, $1-\beta=0.9$, and varying potential number of observed deaths $d$. Dashed curve indicates discrete thresholds that approximate fleming by maintaining $1-\beta\geq 0.9$ while tolerating higher levels of $\alpha$.
Regardless of the approach taken, the low mortality rate in indolent diseases makes rigorous OS monitoring challenging. We agree with fleming that any set of proposed thresholds should be seen as non-binding “informative guidelines” as opposed to black-and-white “rules” for continuing or stopping a trial. At a minimum, the need to monitor OS for safety should encourage stakeholders to contemplate the level of OS data that is likely to be accessible during the course of a clinical trial. Meanwhile, the additional analyses suggested by rodriguez may be helpful for gaining a better understanding of immature OS data: analysis of biologically plausible subgroups, simulation and supplementary analyses to investigate the robustness of results under different assumptions, evaluation of other endpoints, results from other trials, use of real-world data, etc.. IDMCs may also find measures of stability or graphical summaries such as the predicted interval plot useful for understanding potential effect size estimates and associated precision with trial continuation (Betensky2015; Li2009).
@@
  • How should $\theta^{*}$ be defined given potential confounding by intercurrent events? If a substantial number of patients switch to subsequent therapy after tumor progression, then it may be desirable to relax $\theta^{*}$ since subsequent therapies can lead to a smaller OS difference (Korn2011).
@@
Another question that may come up is regarding non-proportional hazards (NPH). Is the proposed framework applicable under NPH? Our opinion is yes, that at least a good part of the framework can be applied if the overall HR is still considered a meaningful treatment effect summary; for example, analysis timing can still be based on the number of deaths $d$ and $\theta^{*}$ can still be used as the threshold for signaling overall OS detriment. However, if it is known a priori that a NPH effect could exist, then operating characteristics should be evaluated carefully under potential patterns of NPH in order to ensure that whatever plan is proposed for OS safety analysis balances risk appropriately over time. Furthermore, supportive analyses such as piecewise exponential and cure rate models can be prospectively planned to help interpret trial results. This is similar to how some sponsors currently handle NPH for efficacy analysis. They conduct event-driven analyses using the logrank test and cox HR, but consider potential NPH when calculating the required number of events. For example, see Checkmate-274 and Keynote-966 (Bajorin2021; Kelley2023).
This paper proposed a supportive decision making framework based on a single criteria $\theta^{*}$. It is worth noting that rodriguez considered having a ‘grey zone’, which, if a trial were to fall into, would prompt further OS follow-up. Doing so could have the effect of ruling out substantial OS detriment while limiting rejection of treatments with marginal OS benefit. While we did not consider such an extension here for simplicity of comparison and presentation, we do find this idea of practical interest worth pursuing in future discussions and research.